Obtain an expression for the potential at a point near an infinitely long charged wire.
1
Expert's answer
2019-05-08T09:33:31-0400

Lets use Gauss's law for electric field.

Fro wire spartial symmetry, choose surface as cylider with wire as axis of symmetry, with radius r and height h.

Charge of wire per meter "\tau", so charge, enclosed in cylinder is "q = \tau \cdot h"

Electric field directed against wire and perpendicular to it. Therefore, flux of electric field via base and top of cylinder equals to 0. Therefore:


"E(r)\cdot 2\pi r h = q/\varepsilon_0 = \tau h /\varepsilon_0"

"E(r) = \frac{\tau}{ \varepsilon_0 2\pi r}"

To obtain an expression for the potential, integrate "E(r)" with opposite sign:


"\phi(r)-\phi(a) = - \int_{a}^{r}\frac{\tau}{ \varepsilon_0 2\pi r}dr = -\frac{\tau}{ 2\pi \varepsilon_0}\ln{\frac{r}{a}}"

If we choose "\phi(a) = 0" , finally


"\phi(r) = -\frac{\tau}{ 2\pi \varepsilon_0}\ln{\frac{r}{a}}"

where r is the distance to wire.